OpenStudy (anonymous):

A right triangle has base (x - 6) units and height (6x - 12) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials.

OpenStudy (radar):



OpenStudy (anonymous):

A. Pythagoras: h^2 = a^2 + b^2 B. Area = (1/2)(base)(height) C. beats me!

OpenStudy (radar):

Part A (6x-12)^2 + (x - 6)^2 (36x^2 - 144x +144) + (x^2 - 12x + 36) Combine like terms and that is Part A.

OpenStudy (radar):

\[37x ^{2}-156x+180\]

OpenStudy (anonymous):

So Is that the answer for A? Or do you have to simplify it more?

OpenStudy (anonymous):

Should that not be the square root of 37x^2-156X+180?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

a. asks for the square b. asks for the area

OpenStudy (radar):

Sorry, I was out. No that is simplified enough, and the problem was asking for the "square" as @douglaswinslowcooper stated in his post above. Part B is asking for the area. The area A of a triangle is A=(1/2)bh where b is base and h is height. Refer to diagram above A=(1/2)(x-6)(6x-12) simplify by taking 1/2 of the height h getting (3x-6) now multiply (x-6)(3x-6) getting: 3x^2 - 24x + 36 sq units (as it is area.)

OpenStudy (radar):

Part C. Just show that the multiplication of two polynomials (x-6)(3x-6) results in a polynomial 3x^2 - 24x + 36 as a result of the distributive process and combing of similar terms.